### v1

# Hierarchical Balance Packing: Towards Efficient Supervised Fine-tuning for Long-Context LLM

Abstract
Training Long-Context Large Language Models (LLMs) is challenging, as hybrid training with long-context and short-context data often leads to workload imbalances. Existing works mainly use data packing to alleviate this issue but fail to consider imbalanced attention computation and wasted communication overhead. This paper proposes Hierarchical Balance Packing (HBP), which designs a novel batch-construction method and training recipe to address those inefficiencies. In particular, the HBP constructs multi-level data packing groups, each optimized with a distinct packing length. It assigns training samples to their optimal groups and configures each group with the most effective settings, including sequential parallelism degree and gradient checkpointing configuration. To effectively utilize multi-level groups of data, we design a dynamic training pipeline specifically tailored to HBP, including curriculum learning, adaptive sequential parallelism, and stable loss. Our extensive experiments demonstrate that our method significantly reduces training time over multiple datasets and open-source models while maintaining strong performance. For the largest DeepSeek-V2 (236B) MOE model, our method speeds up the training by 2.4 with competitive performance.
1 Introduction
Large Language Models (LLMs) (Dubey et al., 2024; Yang et al., 2024; Cai et al., 2024) have achieved state-of-the-art performance in tasks like machine translation, summarization, and code generation. However, many applications demand to process and understand long-context information (Chen et al., 2023a; Peng et al., 2023; Bai et al., 2024a), such as summarizing books, analyzing legal documents, or retaining context in multi-turn conversations. This underscores the necessity for long-context LLMs that can efficiently process long input sequences.
As mentioned in (Dubey et al., 2024), incorporating long context during the Supervised Fine-Tuning (SFT) (Ouyang et al., 2022) stage is highly necessary. On the other hand, general short-context data is also crucial to maintain the model’s general capabilities. However, this hybrid dataset composition introduces significant challenges, primarily in terms of speed and accuracy. For speed, long-context data intensifies training inefficiencies due to imbalanced workloads; For accuracy, long-context data can degrade performance on short-context tasks, affecting the model’s general capabilities. These challenges hinder the efficiency and effectiveness of SFT for long-context LLMs.
The workload imbalance caused by the hybrid of long and short data arises from two main aspects: (1) within mini-batch imbalance (Yao et al., 2024), which is caused by excessive padding from randomized mini-batch construction, and (2) across mini-batch imbalance (Yao et al., 2024), which is caused by uneven computation distribution over data parallel replicas. Existing approaches mainly address this issue using data packing (Bai et al., 2024a; Wang et al., 2024a), which combines variable-length data into fixed-length mini-batches. While data packing helps mitigate workload imbalances, it can also introduce new challenges. First, data packing alters the data distribution, which might affect the models’ performance. Second, the complexity of attention computation for short- and long-context data differs significantly. As illustrated in Figure 1, simply combining them has a high variance of data length, resulting in imbalanced attention computation and workload imbalance. Third, handling long-context data requires sequential parallelism (SP) and collective communication for attention computation (Jacobs et al., 2023; Liu & Abbeel, 2023), which short data do not need. When simply mixed together, short-context data must wait for long-context data, leading to wasted communication time. Larger models, such as DeepSeek-V2 (236B) (DeepSeek-AI, 2024) MOE (Mixture of Experts) models, introduce higher communication overhead due to the increased number of parameters.
To overcome the limitations of data packing, we propose Hierarchical Balance Packing (HBP), an innovative method that proposes multi-level data packing instead of conventional single-level data packing. HBP consists of three key components: (1) what are the optimal packing groups? (2) How to assign training samples to their optimal group? (3) How to train long-context LLM with those data? Firstly, we propose hierarchical group auto-selection to determine the optimal packing-length group set and corresponding configurations, including the packing length, Gradient Checkpointing configuration (Li et al., 2014), and the SP degree (how many partitions the data is divided into). Secondly, we propose balance packing to allocate each sample to the optimal group, aiming to minimize imbalanced attention computation and communication overhead. Thirdly, we adopt alternative training between different packing groups, along with curriculum learning and a stable loss normalizer to stabilize the training process.
We validate the effectiveness of our method through extensive experiments in multiple settings. For example, on datasets Tulu3 (Lambert et al., 2024a) (32K) + Longcite (Zhang et al., 2024) (128K), our approach speeds up 2.4 (57.1 to 23.8 GPU Days) on DeepSeek-V2 (236B) (DeepSeek-AI, 2024); On datasets OpenHerme (Teknium, 2023) (4K) + Longcite (128K), our approach reduces training time from 2.95 to 2.04 GPU Days about 1.45 speeds up on LLama3.1-8B (Dubey et al., 2024). More importantly, our method preserves performance on both short- and long-context datasets while achieving significant efficiency gains. Experiments on various models at different scales like LLama3.1-8B (Dubey et al., 2024), Qwen2.5-32B (Yang et al., 2024), Qwen2.5-72B (Yang et al., 2024), and DeepSeek-V2 (236B) demonstrate consistent improvements, showing the effectiveness and generalizability of our approach.
2 Related Works
2.1 Long-Context LLM
Long-Context Extension. Long-Context Extensions aim to enhance LLMs’ capabilities in handling long contexts. Current research can be broadly categorized into approaches that require fine-tuning and those that operate in a zero-shot manner. Zero-shot approaches often leverage techniques such as prompt compression (Jiang et al., 2023) or specially designed attention mechanisms (Sun et al., 2023; Han et al., 2024). On the other hand, fine-tuning methods primarily focus on extending position encoding, such as RoPE-based approaches (Peng et al., 2023), or utilizing memory-augmented architectures (Packer et al., 2023).
Long-Context Supervised Fine-Tuning. SFT aligns LLMs more effectively with user intent. For Long-Context SFT, research mainly concentrates on generating long-context datasets (Xiong et al., 2023) and establishing corresponding benchmarks (Chen et al., 2023b; Zhang et al., 2024; Bai et al., 2024b). In contrast, our work emphasizes training efficiency and performance. The most relevant work to ours is LongAlign (Bai et al., 2024a), which also focuses on issues related to workload balance and accuracy degradation. They proposed using packing and loss-reweighting to mitigate these issues. However, they failed to recognize the imbalanced attention computation and wasted communicated overhead due to the packing of short- and long-context data, which limits efficiency.
2.2 Data Packing
Data packing (Wang et al., 2024a) is a more practical approach compared to randomly organizing data batches in LLM training. It reduces padding within batches and minimizes idle time across different data-parallel groups. Common packing methods include Random Packing (Contributors, 2023b), Sorted Batching (Kundu et al., 2024), First Fit Shuffle (FFS), First Fit Decrease (FFD), Best Fit Shuffle (BFS), Shortest-Pack-First Histogram-Packing (SPFHP) (Krell et al., 2021), Iterative sampling and filtering (ISF) (Yao et al., 2024). However, all those packing methods operate on a fixed length. In contrast, our method introduces multiple packing groups with varying lengths, which enables more flexible handling of hybrid training involving short- and long-context data.
3 Problem Analysis
In this section, we first define the notations in Table 1 and introduce performance metrics in Section 3.1. We then conduct a preliminary analysis of the commonly used packing methods in Section 3.2 and training strategies in Section 3.3.
| Symbol | Definition |
| token number in one device | |
| number of devices | |
| local batch size in one device | |
| token number of -th sample in | |
| computation complexity of attention | |
| maximal token number across devices | |
| maximal attention computation across devices | |
| total number of training iterations | |
| token number for SP communication in one iteration |
3.1 Measuring Metrics
Dist Balance Ratio (DBR) (Yao et al., 2024) quantifies the computational balance inter-devices based on input length.
Padding Ratio (PR) (Yao et al., 2024) measures the proportion of wasted computations resulting in intra-device from padding based on input length.
Attention Balance Ratio (ABR) is proposed to quantify the imbalance with respect to attention computation for different data inter-devices.
Previous metrics estimate the computational cost of attention based solely on the length of the input with full attention. However, using packing algorithms, attention computation becomes a significant factor in the overall cost. Consider the packing of a sequence as an example, both and have the same total length. However, their actual attention computation differs significantly, with complexities of and , respectively. The Attention Balance Ratio (ABR) is given by:
The cost of attention is proportional to the attention balance ratio, Cost(Attn) ABR.
Communication Ratio (CR) is proposed to measure the additional communication overhead.
Sequence parallelism (SP) is crucial in long-context LLM training. However, this introduces the trade-off of increased communication overhead. The additional communication during hybrid training depends on the input sample length. For instance, an 8k input may not require any SP partitions, while a 32k input might necessitate multiple SP partitions, significantly increasing communication costs.
Average Tokens (Ave-T) is defined as the average number of tokens processed per iteration, serving as a measure of the model’s workload.
A small token count per iteration indicates computational inefficiency. Increasing the batch size appropriately can improve Ave-T throughput and overall training efficiency.
3.2 Packing Analysis
The packing strategy is widely utilized in SFT to improve training efficiency. We first conducted a comprehensive analysis to validate its effectiveness. However, our study also revealed several challenges and limitations associated with this line of approaches.
Importance of Packing. In Table 2, we conduct three batching alternatives, random batching, ISF packing batching (Yao et al., 2024) (comparison between different packing methods is shown in Table 8), sorted batching (Bai et al., 2024a) (ensures that the sequences within each batch have similar lengths) at three sequence length 4K, 32K, 128K using Tulu3 (Lambert et al., 2024a) dataset. Both the sorted and packing strategies significantly reduce DBR and PR, leading to substantial improvements in training speed. However, in longer sequence scenarios such as 128K, the packing strategy has a higher average of tokens (Ave-T) compared to the sorted method, achieving a higher GPU utilization. This underscores the importance of the packing strategy for training on mixed-length datasets.
| Seq Len | Batching | DBR | PR | Ave-T | GPU Days |
| (speed up) | |||||
| 4K | random | 0.540 | 0.416 | 2.4K | 8.0 (1.0) |
| 4K | sorted | 0.001 | 0.001 | 2.4K | 3.3 (2.4) |
| 4K | packing | 0.003 | 0.0 | 4K | 3.1 (2.6) |
| 32K | random | 0.64 | 0.0 | 0.8K | 16.7 (1.0) |
| 32K | sorted | 0 | 0.0 | 0.8K | 10.3 (1.6) |
| 32K | packing | 0.0007 | 0.0 | 32K | 4.4 (3.8) |
| 128K | random | 0.639 | 0.0 | 0.9K | 38.0 (1.0) |
| 128K | sorted | 0 | 0.0 | 0.9K | 33.3 (1.1) |
| 128K | packing | 0.001 | 0.0 | 128K | 5.2 (7.3) |
Limitation of Packing. Although packing can partially address the efficiency issues associated with hybrid training, several limitations remain. Specifically, samples with varying sequence lengths exhibit different complexities in attention computation. Directly mixing these samples can result in a workload imbalance. As illustrated in Table 3, the ABR increases significantly with the growth of sequence length, which indicates a rise in device idle time. Moreover, long sequences necessitate communication for attention computation, while short sequences do not. Directly mixing them can lead to extra communication overhead. As shown in Table 3, the CR reaches 1 when the sequence length increases, indicating that all short-context data are involved in unnecessary communication processes.
| Packing Len | SP | ABR | CR | DBR | PR |
| 4K | 1 | 0.434 | 0 | 0.003 | 0.0 |
| 32K | 4 | 0.616 | 1.0 | 0.001 | 0.0 |
| 128K | 8 | 0.667 | 1.0 | 0.003 | 0.0 |
3.3 Training Strategy Analysis
Given the GPU resources and the data to be trained, we can select from various training strategies, provided that the VRAM requirements are satisfied. The main factors to consider are the degree of SP and the configuration of Gradient Checkpointing (GC), which is the number of layers where GC is enabled. If the SP degree is small, the VRAM demand is high, which forces an increase in the number of GC layers and leads to excessive additional computation. On the other hand, if the SP degree is large, although the VRAM demand is reduced, it introduces additional communication overhead. Therefore, there is a trade-off between the SP degree and GC configuration. Moreover, the optimal strategy varies for different sequence lengths. Table 4 shows that the optimal strategies for 32K, 64K, and 128K are different.
| Seq Len | SP | GC Layer | Memory | Iter Time |
| 32K | 2 | 28 | 77G | 4.45s |
| 32K | 4 | 23 | 78G | 4.35s |
| 32K | 8 | 8 | 78G | 4.12s |
| 64K | 2 | 32 | OOM | - |
| 64K | 4 | 28 | 78G | 6.3s |
| 64K | 8 | 24 | 79G | 6.2s |
| 128K | 4 | 32 | OOM | - |
| 128K | 8 | 29 | 78G | 10.2s |
| 128K | 16 | 23 | 79G | 10.5s |
4 Hierarchical Balance Packing
In this section, we first introduce how to determine the optimal packing length groups in Section 4.1. Then, we present how to fit each sample to the best packing group in Section 4.2. At last, we design a tailored dynamic training pipeline for HBP in Section 4.3. The overall framework is illustrated in Figure 2.
4.1 Hierarchical Groups Auto-Selection
To determine the optimal packing length groups, we design a profile-based auto-selection algorithm as described in Algorithm 1. It operates in two stages: (1) find the best training strategy for predefining possible sequence length set (e.g., 8K, 16K, 32K, 64K, 128K) based on naive packing. (2) deriving the final packing groups by optimizing communication overhead.
Stage 1: Find the best training strategy. The algorithm begins by initializing two empty lists, and , to store the profiling times and the corresponding strategies, respectively. For each possible input length , we compute the optimal SP degree and GC configuration with the FindBestSpCkpt function, which achieves the optimal trade-off between SP degree and GC configuration.
Specifically, given a packing length , we iterate all possible degrees and GC configurations and profile their iteration time. The best combination is then found by a binary search or greedy method. More details on the implementation can be found in Appendix B. Once all input lengths are processed, the algorithm selects the best strategy by identifying the index that minimizes the profiling times in . The optimal strategy and input length are:
Additionally, the maximum input length and its corresponding strategy are also retained.
Stage 2: Optimize packing groups for communication.
After calculating the optimal and the corresponding , we can optimize it further. The can an individual sample to be split across different SP processes, leading to additional communication overhead during attention computation. To mitigate this, we propose a method that ensures smaller samples remain intact during SP training. When and , instead of directly packing data into 32K sequences, we first partition into smaller chunks of . Data with lengths smaller than 8K are packed into these 8K chunks first, which are then combined to form 32K sequences. By structuring the data this way, SP training can proceed without incurring any communication overhead from attention computation.
For , the smallest packing unit is derived based on its optimal SP degree. Similarly, the smallest packing unit for is also considered. If is smaller than , it will be merged into the range. Finally, the hierarchical packing groups is obtained.
4.2 Balance Packing
After obtaining the optimal hierarchical pack groups, we distribute the dataset samples into different groups while ensuring that the metrics outlined in Section 3.1 (DBR, PR, ABR, and CR) are well-optimized, of which conventional packing struggles. We first divide the entire dataset into sub-datasets based on . For each , the following steps are executed:
(1) Packing: We pack the data in to length . This ensures low PR and DBR. Note that arbitrary packing methods are feasible.
(2) GreedyFill: We use the remaining unpacked data to perform a greedy fill since larger groups find it difficult to fill the packed data with only their own data.
(3) Sorting and Batching: We sort elements in based on attention complexity and construct mini-batches according to global token number requirements.
The procedure of balance packing is illustrated in Algorithm 2. Since our approach inherently involves multiple levels, i.e., hierarchical packing groups, it automatically separates short- and long-context data, avoiding wasted communication overhead and imbalanced attention computation and reducing CR and ABR significantly. We also achieve extremely low PR and DBR at multiple levels, thanks to GreedyFill. More Details about GroupData, GreedyFill, and Sorting are shown in Appendix C.
| Model | General Tasks | Long Tasks | GPU Days | ||||||||
| (Type) | AVE | MMLU | BBH | IFEval | Math | GSM8k | HumanEval | Ruler (32K128K) | LongBench | LongCite | (speed up) |
| LLama3.1-8B | |||||||||||
| LongAlign-packing | 56.6 | 44.5 | 65.5 | 67.8 | 30.7 | 80.6 | 62.2 | 84.5 57.5 | 46.7 | 67.8 | 7.41 (0.7) |
| LongAlign-sorted | 57.6 | 62.7 | 65.4 | 67.8 | 32.8 | 82.2 | 61.6 | 85.8 59.9 | 46.5 | 64.0 | 33.3 (0.16) |
| ISF | 56.0 | 54.5 | 65.3 | 70.4 | 33.7 | 81.7 | 62.8 | 85.0 67.4 | 44.0 | 71.6 | 5.22 (1.0) |
| HBP | 58.2 | 63.0 | 67.2 | 67.7 | 33.0 | 81.9 | 63.4 | 85.6 70.8 | 43.1 | 71.5 | 3.73 (1.4) |
| Qwen2.5-32B | |||||||||||
| ISF | 73.5 | 74.8 | 83.5 | 75.6 | 56.5 | 93.7 | 86.6 | 88.2 59.3 | 51.0 | 60.2 | 21.3 (1.0) |
| HBP | 76.2 | 76.6 | 83.6 | 76.0 | 57.1 | 94.4 | 84.2 | 88.3 59.0 | 51.9 | 61.7 | 16.0 (1.33) |
| LLama3.1-70B | |||||||||||
| ISF | 72.1 | 78.9 | 83.0 | 77.6 | 44.1 | 85.3 | 76.8 | 91.8 57.1 | 50.4 | 72,7 | 44.4 (1.0) |
| HBP | 74.2 | 81.5 | 83.1 | 76.2 | 48.3 | 93.3 | 77.4 | 93.4 57.5 | 52.2 | 75.3 | 31.1 (1.42) |
| DeepSeek-V2 (236B) | |||||||||||
| ISF | 71.8 | 76.8 | 84.0 | 71.1 | 41.3 | 89.0 | 78.6 | 86.6 - | 47.1 | - | 57.1 (1.0x) |
| HBP | 72.0 | 76.5 | 83.1 | 72.6 | 41.4 | 89.9 | 78.1 | 87.3 - | 50.3 | - | 23.8 (2.4) |
4.3 Dynamic Training Pipeline
Since HBP involves multi-level inputs, it is essential to design a dynamic training pipeline, enabling hot switching of different packing groups. It incorporates an adaptive sequential parallel, a curriculum learning strategy, and a stable loss normalizer.
Adaptive Sequential Parallel: Each packing group is equipped with an optimal training strategy . We use alternative training between packing groups with the best SP degree and GC configuration.
Curriculum Learning Strategy: Training on long-context tasks presents challenges because initiating training without instructional capabilities can result in significant fluctuations in the training loss, as illustrated in Figure 3. Thanks to our inherent hierarchical structure, it is much easier to adopt a curriculum learning strategy that starts with general short-context data during the early stages of training. As the training progresses, we shift to a hybrid approach that combines both general short- and long-context data.
Stable Loss Normalizer: The training stability introduced by data packing is an important problem as it impacts the data distribution. Previous work (Bai et al., 2024a) has analyzed loss calculation, identifying two primary loss normalizers Token Mean and Sample Mean:
where represents the batch size, and denotes the number of loss tokens in the current batch. They argued that instability lies in the inconsistent loss of normalization. (Lambert et al., 2024b) also proposes sum loss without normalization to mitigate the effect of norm discrepancies. However, the sum loss introduces a trade-off: as the sequence length increases, the gradient values escalate disproportionately (1e+5), as shown in the left part of Figure 4.
To address the above issues, we empirically find (Average Token) of the global batch size could served as a Stable Loss Normalizer:
where represents the number of loss tokens that need to be calculated for the current batch .
5 Experiments
5.1 Experimental Setup
We use large-scale datasets: Tulu3 (32K) (Lambert et al., 2024a) (general task), and LongCite (128K) (Zhang et al., 2024) (long context task). These datasets have proven effective in previous research, significantly enhancing model performance across knowledge, reasoning, mathematics, coding, and instruction-following. Moreover, they improve the model’s ability to handle varying lengths, from 0.1K to 128K tokens.
Implementation Details. We conduct experiments with the following models: LLaMA 3.1 (Dubey et al., 2024), Qwen-2.5 (Yang et al., 2024), and DeepSeek-V2 (236B). Most models are trained on 32x H100 80GB GPUs using the DeepSpeed (Rajbhandari et al., 2020), while DeepSeek-V2 (236B) is trained with the Megatron-LM (Shoeybi et al., 2019) with 256x H100 80G GPUs. We conducted ablation experiments using the LLaMA3.1-8B model. For the Longsite dataset, approximately 2k samples are uniformly sampled. The loss normalizer for baselines without special instructions is Token-Mean, while HBP uses Ave-Token. GPU days are the evaluation metric to estimate the total training time.
| Dataset | Pack Len | AVE | LongBench | Ruler-128K |
| Tulu3 | 32K | 57.5 | 43.0 | 52.5 |
| Longcite | 128K | 18.8 | 16.7 | 68.5 |
| Tulu3 + Longcite(8K) | 32K | 58.6 | 43.2 | 62.1 |
| Tulu3 + Longcite | 128K | 56.0 | 44.0 | 67.5 |
| Model | Hierarchical | Balance | ABR | CR | AVE | GPU Days (speed up) |
| LLaMA3.1-8B | 0.506 | 1.0 | 56.0 | 5.22 (1.0) | ||
| LLaMA3.1-8B | ✓ | 0.288 | 0.173 | 56.4 | 4.51 (1.2) | |
| LLaMA3.1-8B | ✓ | ✓ | 0.002 | 0.173 | 56.6 | 3.73 (1.40) |
| LLaMA3.1-70B | 0.506 | 1.0 | 72.2 | 44.4 (1.0) | ||
| LLaMA3.1-70B | ✓ | 0.288 | 0.173 | 72.8 | 33.3 (1.25) | |
| LLaMA3.1-70B | ✓ | ✓ | 0.002 | 0.173 | 72.2 | 31.1 (1.43) |
Evaluation. We conducted a comprehensive evaluation of the LLM’s performance using OpenCompass (Contributors, 2023a). For general tasks, several benchmark datasets were assessed, including MMLU (Hendrycks et al., 2021), MMLU PRO (Wang et al., 2024b), CMMLU (Li et al., 2023), BBH (Suzgun et al., 2022), Math (Saxton & Kohli, 2019), GPQA Diamond (Rein et al., 2024), GSM8K (Cobbe et al., 2021), HellaSwag (Zellers et al., 2019), MathBench (Liu et al., 2024), HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), IFEval (Zhou et al., 2023), and Drop (Dua et al., 2019). For long-context tasks, the evaluation including Ruler (Hsieh et al., 2024), NeedleBench (Li et al., 2024), LongBench (Bai et al., 2023), and Longcite.
5.2 Main Results
In Table 5, we compare our method HBP with Long-Align (Bai et al., 2024a) and packing method ISF (Yao et al., 2024). Our method achieves significant improvements in speed, outperforming Long-Align-packing, Long-Align-sorted, ISF by 2.0, 9.0, and 1.4. We also notice that LongAlign improves the average general tasks to some extent; it sacrifices performance in long tasks e.g., Ruler-128K. In contrast, our method maintains strong performance both in general short tasks and long tasks. The improvements are consistent across a wide range of model sizes, from 8B to 236B parameters. Notably, for the largest MOE model, DeepSeek-V2 (236B), our method achieves an impressive 2.4 training speed-up, reducing training time from 57.1 to 23.8 GPU days. Full results are shown in Appendix A.
5.3 Packing Strategy Results
Table 8 and Table 9 illustrate the complexity of different packing strategies and their corresponding final training times. It also shows that the training indicators Data Balance Ratio (DBR), Padding Ratio (PR), and Attention Balance Ratio (ABR) are strongly correlated with the training time, emphasizing the effectiveness of these indicators. Based on the computational complexity of packing strategies and their training time and accuracy performance, we ultimately selected ISF as our naive packing baseline.
| Packing Strategy | Complexity | DBR | ABR | PR | Ave | LongBench | GPU Days (speed up) |
| No | - | 0.64 | 0.744 | 0 | 57.6 | 42. | 16.7 (1.00) |
| Random | O(N) | 0 | 0.620 | 0.05 | 57.6 | 42.7 | 4.44 (3.76) |
| ISF | C*O(N+M) | 0 | 0.616 | 0.002 | 58.1 | 43.8 | 4.00 (4.17) |
| FFS | O(NM) | 0 | 0.618 | 0.001 | 8.3 | 42.1 | 4.02 (4.15) |
| FFD | O(NM) | 0 | 0.693 | 0.001 | 58.0 | 43.1 | 4.51 (3.70) |
| BFS | O(NM) | 0 | 0.618 | 0.001 | 57.5 | 42.1 | 4.11 (4.06) |
| SPFHP | O(N+S2) | 0 | 0.695 | 0.001 | 58.6 | 40.6 | 4.24 (3.94) |
| Packing Strategy | Complexity | DBR | ABR | PR | Ave | LongBench | GPU Days (speed up) |
| - | - | 0 | 0.878 | 0.676 | 48.2 | 46.4 | 11.4 (1.0) |
| random | O(N) | 0 | 0.648 | 0.089 | 51.44 | 47.8 | 3.64 (3.1) |
| ISF | C*O(N+M) | 0 | 0.64 | 0.022 | 50.9 | 47.8 | 3.28 (3.5) |
| FFS | O(NM) | 0 | 0.638 | 0.022 | 52.4 | 48.4 | 3.29 (3.5) |
| FFD | O(NM) | 0 | 0.71 | 0.022 | 50.6 | 48.4 | 3.45 (3.3) |
| BFS | O(NM) | 0 | 0.64 | 0.022 | 52.3 | 47.7 | 3.33 (3.4) |
| SPFHP | O(N+S2) | 0 | 0.71 | 0.029 | 52.0 | 47.8 | 3.47 (3.3) |
5.4 Ablation Results
Importance of Hybrid Training. Table 6 shows that both short-context and long-context data are important to maintain models’ capability in general tasks and long tasks. Lacking of long-context data (row 1) significantly impacts long-text capabilities. Lacking short-context data (row 2) sacrifices general abilities. These claims are further validated by only including partial short data in Longcite (row 3).
Components of HBP. In Table 7, we show that the Attention Balance Ratio (ABR) and Communication Ratio (CR) can be reduced significantly with hierarchical packing. In particular, ABR drops from 0.506 to 0.288, and the CR drops from 1.0 to 0.173. By batching data with a similar complexity of attention computation, we have much more balanced mini-batches with low ABR, from 0.288 to 0.002. Overall, we achieve a 1.4 speedup. Similar results can be observed in the larger LLaMA-3.1-70B model.
Curriculum Learning. In Table 10, we present the impact of curriculum learning on HBP. It is evident that starting with short general tasks and then transitioning to a mix of short- and long-context tasks is more beneficial for model training and convergence. We also designed a similar curriculum learning mechanism for the naive ISF baseline using a sophisticated sampling strategy, which shows improvements. This demonstrates that curriculum learning is a generalizable strategy for long-context SFT. Notably, since HBP naturally separates short and long contexts into different groups, employing curriculum learning becomes more straightforward and convenient.
| Model | CL | AVE | LongBench |
| LLama3.1-8B-HBP | 56.6 | 41.6 | |
| LLama3.1-8B-HBP | ✓ | 58.2 | 43.2 |
| LLama3.1-70B-HBP | 72.2 | 51.5 | |
| LLama3.1-70B-HBP | ✓ | 74.1 | 52.2 |
| LLama3.1-8B-ISF | 56.0 | 44.0 | |
| LLama3.1-8B-ISF | ✓ | 57.4 | 43.4 |
Stable Loss Normalizer. We compared several loss normalization methods by training models using different normalizers while keeping all other training configurations consistent. As shown in Table 11, the Ave-Token loss normalizer achieved the highest performance in both general tasks Average (AVE) 58.2 and long context tasks LongBench 43.1, Ruler-128K 70.8, and LongSite 71.5.
| Loss Normalizer | AVE | LongBench | Ruler-128K | Longcite |
| Sum | 56.7 | 42.5 | 65.2 | 70.5 |
| Sample-Mean | 55.5 | 42.9 | 46.1 | 70.3 |
| Token-Mean | 56.6 | 41.6 | 67.5 | 70.6 |
| Ave-Token | 58.2 | 43.1 | 70.8 | 71.5 |
5.5 Importance of Hierarchical Groups Auto-Selection
Table 12 presents an example of 32K token-length training, showcasing various SP degrees and GC configurations. The second row highlights the minimal configuration that satisfies memory constraints. While this configuration adheres to memory requirements, it fails to achieve optimal performance. In contrast, the fourth row illustrates a more balanced and effective configuration with , achieving the fastest speed. These experiments demonstrate that, given a specific packing length, there are significant performance differences among various SP degrees and GC configurations. Table 13 presents the training speed of the optimal SP degrees and GC configurations for various sequence lengths . This indicates that the optimal training strategy is distinct for different packing groups. The evidence above emphasizes the importance of searching for the best groups and their corresponding training strategies, i.e., Auto-Group Selection.
| SP | GC Layer | AVE | Memory | Iter Time | GPU Days |
| 1 | 32 | - | OOM | - | - |
| 2 | 28 | 57.5 | 78G | 3.01 s | 3.73 |
| 4 | 23 | 57.8 | 78G | 2.95 s | 3.37 |
| 8 | 8 | 57.6 | 77G | 2.82 s | 3.04 |
| 16 | 0 | 56.6 | 76G | 3.60 s | 3.93 |
| Packing Groups | SP | GC Layer | Memory | Iter Time |
| 8k | 2 | 8 | 77 G | 2.69 s |
| 16k | 1 | 28 | 76 G | 2.65 s |
| 32k | 8 | 8 | 76 G | 2.83 s |
| 64k | 4 | 28 | 78 G | 3.01 s |
| 128k | 8 | 28 | 79 G | 3.05 s |
| Hierarchical | Balance | ABR | AVE | LongBench | Ruler-32K | Longsite | GPU Days (speed up) |
| 0.513 | 52.1 | 47.2 | 87.7 | 72.1 | 2.95 (1.0) | ||
| ✓ | 0.269 | 53 | 47.1 | 89.1 | 72.0 | 2.33 (1.27) | |
| ✓ | ✓ | 0.004 | 52.4 | 47.1 | 88.3 | 72.3 | 2.04 (1.44) |
| Model | Dataset | AVe | LongBench | LongWriter | GPU Days (speed up) |
| Llama3.1-8B-ISF | Tulu3 + LongWriter | 57.3 | 40.9 | 67.0 | 4.4 (1.0) |
| Llama3.1-8B-HBP | Tulu3 + LongWriter | 57.8 | 42.7 | 66.4 | 3.5 (1.26) |
5.6 Dataset Generalization
5.6.1 OpenHermes
5.6.2 LongWriter
6 Conclusion and Limitations
In this paper, we proposed Hierarchical Balance Packing (HBP), a novel strategy to address workload imbalances in long-context LLM training through multi-level data packing and a dynamic training pipeline. Due to constraints in computational resources and open-source training datasets, we have not tested on longer contexts, such as 256K or 512K. Additionally, we have not validated HBP on other post-training tasks, such as RLHF or DPO. We leave these explorations to future work.
Appendix A Full Results
Tables 16 and 17 present our complete results for general tasks and long-context tasks, respectively. These results collectively validate the effectiveness of our HBP method.
| Model | MMLU | BBH | IFEval | Math | GSM8k | HumanEval | mmlu_pro | cmmlu | GPQA | Drop | MBPP | hellaswag | mathbench-a | mathbench-t | AVE | GPU Days (speed up) |
| LLama3.1-8B | ||||||||||||||||
| LongAlign-packing | 44.5 | 65.5 | 67.8 | 30.7 | 80.6 | 62.2 | 26.5 | 48.4 | 28.8 | 71.5 | 63.8 | 80.5 | 45.6 | 75.6 | 56.6 | 7.4 (0.7) |
| LongAlign-sorted | 62.7 | 65.4 | 67.8 | 32.8 | 82.2 | 61.6 | 38.3 | 43.6 | 29.3 | 65.1 | 63.0 | 69.6 | 50.3 | 73.9 | 57.6 | 33.3 (0.16) |
| ISF | 54.5 | 65.3 | 70.4 | 33.7 | 81.7 | 62.8 | 34.6 | 39.8 | 24.2 | 65.4 | 61.8 | 67.9 | 46.7 | 74.7 | 56.0 | 5.22 (1.0) |
| HBP | 63.0 | 67.2 | 67.7 | 33.0 | 81.9 | 63.4 | 38.4 | 43.4 | 27.8 | 68.7 | 65.0 | 68.7 | 50.1 | 76.4 | 58.2 | 3.73 (1.4) |
| Qwen2.5-32B | ||||||||||||||||
| ISF | 74.8 | 83.5 | 75.6 | 56.5 | 93.7 | 86.6 | 59.6 | 77.6 | 37.9 | 82.7 | 80.1 | 93.5 | 64.1 | 63.5 | 73.5 | 21.33 (1.0) |
| HBP | 76.6 | 83.6 | 76.0 | 57.1 | 94.4 | 84.2 | 59.2 | 79.5 | 41.4 | 83.5 | 80.9 | 93.5 | 70.1 | 86.2 | 76.2 | 16.00 (1.33) |
| LLaMA3.1-70B | ||||||||||||||||
| ISF | 78.9 | 83.0 | 77.6 | 44.1 | 85.3 | 76.8 | 55.0 | 66.9 | 41.4 | 81.9 | 76.6 | 89.3 | 65.1 | 88.0 | 72.1 | 44.40 (1.0) |
| HBP | 81.5 | 83.1 | 76.2 | 48.3 | 93.3 | 77.4 | 60.2 | 66.3 | 43.9 | 84.1 | 78.6 | 89.0 | 67.9 | 88.4 | 74.2 | 31.10 (1.42) |
| Qwen2.5-72B | ||||||||||||||||
| ISF | 83.9 | 86.0 | 79.3 | 57.2 | 94.6 | 85.9 | 66.2 | 84.7 | 45.4 | 84.6 | 84.8 | 93.7 | 74.1 | 95.0 | 79.6 | 47.10 (1.0) |
| HBP | 84.2 | 85.8 | 79.7 | 56.2 | 94.7 | 85.0 | 65.9 | 84.9 | 50.5 | 84.9 | 86.4 | 93.9 | 72.8 | 95.1 | 79.5 | 33.70 (1.40) |
| DeepSeek-V2 (236B) | ||||||||||||||||
| ISF | 76.8 | 84.0 | 71.1 | 41.3 | 89.0 | 78.6 | 54.2 | 76.9 | 37.9 | 77.2 | 76.7 | 90.2 | 68.4 | 92.3 | 71.8 | 57.10 (1.0) |
| HBP | 76.5 | 83.1 | 72.6 | 41.4 | 89.9 | 78.1 | 55.5 | 73.3 | 36.4 | 78.5 | 77.4 | 90.2 | 70.1 | 92.5 | 72.0 | 23.80 (2.4) |
| Model | Ruler (32K | 128K) | NeedleBench (32K | 128K) | LongBench | Longcite | GPU Days (speed up) |
| LLama3.1-8B | |||||
| LongAlign-packing | 84.5 | 57.5 | 87.9 | 85.0 | 46.7 | 67.8 | 7.4 (0.7) |
| LongAlign-sorted | 85.6 | 60.0 | 92.4 | 88.9 | 46.6 | 64.0 | 33.3 (0.16) |
| ISF | 85.0 | 67.4 | 92.1 | 90.1 | 44.5 | 71.6 | 5.22 (1.0) |
| HBP | 85.6 | 70.8 | 91.8 | 90.0 | 43.2 | 71.5 | 3.73 (1.4) |
| Qwen2.5-32B | |||||
| ISF | 88.2 | 59.3 | 94.5 | 84.6 | 51.0 | 60.2 | 21.33 (1.0) |
| HBP | 88.3 | 59.0 | 96.0 | 88.9 | 51.9 | 61.7 | 16.00 (1.33) |
| LLaMA3.1-70B | |||||
| ISF | 91.8 | 57.1 | 95.5 | 92.6 | 50.4 | 72.7 | 44.4 (1.0) |
| HBP | 93.4 | 57.5 | 95.2 | 92.4 | 52.2 | 75.3 | 31.1 (1.42) |
| Qwen2.5-72B | |||||
| ISF | 92.6 | 58.5 | 94.5 | 90.8 | 50.0 | 64.3 | 47.1 (1.0) |
| HBP | 92.8 | 58.2 | 95.2 | 91.4 | 51.7 | 64.7 | 33.7 (1.40) |
| DeepSeek-V2 (236B) | |||||
| ISF | 86.6 | - | 96.0 | - | 47.1 | - | 57.1 (1.0) |
| HBP | 87.3 | - | 95.9 | - | 50.3 | - | 23.8 (2.4) |
Appendix B Details of Hierarchical Groups Auto-Selection
B.1 FindBestSpCkpt
Algorithm 3 determines the optimal gradient checkpointing strategy by evaluating all possible strategies.
Initialization: Start with empty lists and for profiling times and configurations, respectively.
Iterate Over Strategies: For each strategy :
Compute the best gradient checkpointing configuration () using the GreedyProfileCkpt function.
Use ProfileTime to profile the execution time for the model with the given configuration and append it to .
Find Optimal Strategy: Identify the index of the minimum profiling time in using .
Return Best Configuration: Return the SP degree and corresponding gradient checkpointing configuration .
B.2 GreedyProfileCkpt
Algorithm 4 estimates the number of gradient checkpointing layers required for a given and strategy.
Initialization: Obtain the configurations using the minimum and maximum number of gradient checkpointing layers (based on empirical observations):
Memory Profiling: Profile the remaining memory for () and ().
Memory Slope Calculation: Compute the average memory slope () as:
Checkpointing Layer Estimation: Estimate the required number of checkpoints () as:
Appendix C Details of Balance Packing
C.1 GroupData
Given a data set and predefined hierarchical lengths , we evaluate the length of each data set to determine the interval to which it belongs, assigning it to the corresponding group . The detailed procedure is outlined in Algorithm 8.
C.2 GreedyFill
For a given packing group with the corresponding length , we iterate through the smaller dataset partitions and greedily fill the group within the current . The detailed procedure is illustrated in Algorithm 11.
C.3 Attention Balance Sort Function
First, we calculate the attention complexity for all data within the given packing group . Then, we sort the elements in based on their attention complexity and construct mini-batches according to the global token number requirements, as shown in Algorithm 7.