### v3

# Hierarchical Balance Packing: Towards Efficient Supervised Fine-tuning for Long-Context LLM

Abstract
Training Long-Context Large Language Models (LLMs) is challenging, as hybrid training with long-context and short-context data often leads to workload imbalances. Existing works mainly use data packing to alleviate this issue, but fail to consider imbalanced attention computation and wasted communication overhead. This paper proposes Hierarchical Balance Packing (HBP), which designs a novel batch-construction method and training recipe to address those inefficiencies. In particular, the HBP constructs multi-level data packing groups, each optimized with a distinct packing length. It assigns training samples to their optimal groups and configures each group with the most effective settings, including sequential parallelism degree and gradient checkpointing configuration. To effectively utilize multi-level groups of data, we design a dynamic training pipeline specifically tailored to HBP, including curriculum learning, adaptive sequential parallelism, and stable loss. Our extensive experiments demonstrate that our method significantly reduces training time over multiple datasets and open-source models while maintaining strong performance. For the largest DeepSeek-V2 (236B) MoE model, our method speeds up the training by 2.4 with competitive performance. Codes will be released at https://github.com/ModelTC/HBP.
1 Introduction
Large Language Models (LLMs) [1, 2, 3] have achieved state-of-the-art performance in tasks like machine translation, summarization, and code generation. However, many applications demand to process and understand long-context information [4, 5, 6], such as summarizing books, analyzing legal documents, or retaining context in multi-turn conversations. This underscores the necessity for long-context LLMs that can efficiently process long input sequences.
As mentioned in [1], incorporating long context during the Supervised Fine-Tuning (SFT) [7] stage is highly necessary. On the other hand, general short-context data is also crucial to maintain the model’s general capabilities. However, this hybrid dataset composition introduces significant challenges, primarily in terms of speed and accuracy. For speed, long-context data intensifies training inefficiencies due to imbalanced workloads; For accuracy, long-context data can degrade performance on short-context tasks, affecting the model’s general capabilities. These challenges hinder the efficiency and effectiveness of SFT for long-context LLMs.
The workload imbalance caused by the hybrid of long and short data arises from two main aspects: (1) within mini-batch imbalance [8], which is caused by excessive padding from randomized mini-batch construction, and (2) across mini-batch imbalance [8], which is caused by uneven computation distribution over data parallel replicas. Existing approaches mainly address this issue using data packing [6, 9], which combines variable-length data into fixed-length mini-batches. While data packing helps mitigate workload imbalances, it can also introduce new challenges. First, data packing alters the data distribution, which might affect the models’ performance. Second, the complexity of attention computation for short- and long-context data differs significantly. The attention complexity of the packed samples is the sum of the attention complexities of all samples within each packing group. As illustrated in Figure 1, naive packing leads to high variance in attention complexity by directly mixing long and short samples, which causes imbalanced attention computation and workload imbalance. Third, handling long-context data requires sequential parallelism (SP) and collective communication for attention computation [10, 11], which short data do not need. When mixed, short-context data also requires sequence parallel communication, leading to wasted communication time. Larger models, such as DeepSeek-V2 (236B) [12] MoE (Mixture of Experts) models, introduce higher communication overhead due to the increased number of parameters.
To overcome the limitations of data packing, we propose Hierarchical Balance Packing (HBP), an innovative method that proposes multi-level data packing instead of conventional single-level data packing. HBP consists of three key components: (1) What are the optimal packing groups? (2) How to assign training samples to their optimal group? (3) How can a long-context LLM be trained with that data? Firstly, we propose hierarchical group auto-selection to determine the optimal packing-length group set and corresponding configurations, including the packing length, Gradient Checkpointing configuration [13], and the SP degree (how many partitions the data is divided into). Secondly, we propose balanced packing to allocate each sample to the optimal group, aiming to minimize imbalanced attention computation and communication overhead. Thirdly, we adopt alternative training between different packing groups, along with curriculum learning and a stable loss normalizer to stabilize the training process.
We validate the effectiveness of our method through extensive experiments in multiple settings. For example, on datasets Tulu3 [14] (32K) + Longcite [15] (128K), our approach speeds up 2.4 (57.1 to 23.8 GPU Days) on DeepSeek-V2 (236B) [12]; On datasets OpenHerme [16] (4K) + Longcite (128K), our approach reduces training time from 2.95 to 2.04 GPU Days about 1.45 speeds up on LLama3.1-8B [1]. More importantly, our method preserves performance on both short- and long-context datasets while achieving significant efficiency gains. Experiments on various models at different scales like LLama3.1-8B [1], Qwen2.5-32B [2], Qwen2.5-72B [2], and DeepSeek-V2 (236B) demonstrate consistent improvements, showing the effectiveness and generalizability of HBP.
2 Related Works
2.1 Long-Context LLM
Long-Context Extension. Long-Context Extensions aim to enhance LLMs’ capabilities in handling long contexts. Current research can be broadly categorized into approaches that require fine-tuning and those that operate in a zero-shot manner. Zero-shot approaches often leverage techniques such as prompt compression [17] or specially designed attention mechanisms [18, 19]. On the other hand, fine-tuning methods primarily focus on extending position encoding, such as RoPE-based approaches [5], or utilizing memory-augmented architectures [20].
Long-Context Supervised Fine-Tuning. For Long-Context SFT, research mainly concentrates on generating long-context datasets [21] and establishing corresponding benchmarks [22, 15, 23]. LongAlign [6] also focuses on workload balance and accuracy degradation issues. LongAlign proposed using packing and loss-reweighting to mitigate these issues. However, they failed to recognize the imbalanced attention computation and wasted communication overhead due to the packing of short- and long-context data. FlexSP [24] dynamically organizes training samples into different data groups and uses flexible sequence parallelism to enable hybrid training. Its online data organization introduces significant overhead (5-15 seconds) each iteration and fails to handle attention computation complexity across data-parallel (DP) groups. In contrast, our method introduces only negligible overhead and achieves a better computation balance of attention.
2.2 Data Packing
Data packing [9] is a more practical approach compared to randomly organizing data batches in LLM training. It reduces padding within batches and minimizes idle time across different data-parallel groups. Common packing methods include Random Packing [25], Sorted Batching [6], First Fit Shuffle (FFS), First Fit Decrease (FFD), Best Fit Shuffle (BFS), Shortest-Pack-First Histogram-Packing (SPFHP) [26], Iterative sampling and filtering (ISF) [8]. However, those packing methods operate on a fixed length. HBP operates data globally by introducing multiple packing groups of varying lengths, enabling more flexible and efficient handling of hybrid training with both short- and long-context data.
3 Problem Analysis
In this section, we first define the notations in Table 2 and introduce performance metrics in Section 3.1. We then conduct a preliminary analysis of the commonly used packing methods in Section 3.2 and training strategies in Section 3.3.
| Symbol | Definition |
| token number in one device | |
| number of devices | |
| local batch size in one device | |
| token number of -th sample in | |
| computation complexity of attention | |
| maximal token number across devices | |
| maximal attention computation across devices | |
| total number of training iterations | |
| token number for SP communication in one iteration |
| Packing Len | SP | ABR | CR | DBR | PR |
| 4K | 1 | 0.343 | 0 | 0.003 | 0.0 |
| 32K | 4 | 0.456 | 1.0 | 0.001 | 0.0 |
| 128K | 8 | 0.506 | 1.0 | 0.003 | 0.0 |
3.1 Measuring Metrics
Dist Balance Ratio (DBR) [8] quantifies the computational balance inter-devices based on length. A lower DBR 1 indicates more balanced workload distribution across different devices.
Padding Ratio (PR) [8] measures the proportion of wasted computations resulting from intra-device padding based on input length. A lower PR 1 indicates fewer padding tokens.
Attention Balance Ratio (ABR) is proposed to quantify the imbalance in attention computation for different data inter-devices. Previous metrics estimate the computational cost of attention based solely on the length of the input with full attention. However, using packing algorithms, attention computation becomes a significant factor in the overall cost. Consider the packing of a sequence as an example, both and have the same total length. However, their actual attention computation differs significantly, with complexities of and , respectively. The Attention Balance Ratio (ABR) and example are given by 2. A lower ABR indicates a more balanced attention computation between different DP groups, resulting in less idle time.
Communication Ratio (CR) is proposed to measure the overhead of SP communication. in CR 3 represents the total number of tokens that require SP communication in the current iteration, while denotes the total number of tokens in the current iteration. A lower CR 3 indicates that fewer tokens require SP communication, resulting in reduced communication overhead.
Average Tokens (Ave-T) is defined as the average number of tokens processed per iteration, serving as a measure of the model’s workload. A larger Ave-T 3 indicates higher training efficiency.
3.2 Packing Analysis
Importance of Packing. In Table 4, we conduct three batching alternatives, random batching, ISF packing batching [8] (comparison between different packing methods is shown in Appendix A), sorted batching [6] (ensures that the sequences within each batch have similar lengths) at three sequence length 4K, 32K, 128K using Tulu3 [14] dataset. Both sorting and packing reduce DBR and PR, speeding up training. On longer sequences like 128K, packing achieves higher Ave-T than sorting, improving GPU utilization and benefiting mixed-length datasets.
Limitation of Packing. Although packing can partially address the efficiency issues associated with hybrid training, several limitations remain. Imbalance Problem of Attention Complexity: Packing samples with varying sequence lengths leads to differing attention computation complexities. Directly mixing these samples can cause imbalanced workloads and inefficient resource utilization. As illustrated in Table 2, the ABR increases significantly with the sequence length growth, indicating a rise in device idle time. More details are shown in Appendix D. SP Communication Overhead: long sequences require communication for attention computation, whereas short sequences do not. Directly mixing them can lead to extra communication overhead. As shown in Table 2, the CR reaches 1 when the sequence length increases, indicating that all short-context data are involved in unnecessary communication.
| Seq Len | Batching | DBR | PR | Ave-T | GPU Days |
| (speed up) | |||||
| 4K | random | 0.540 | 0.416 | 2.4K | 8.0 (1.0) |
| 4K | sorted | 0.001 | 0.001 | 2.4K | 3.3 (2.4) |
[…]
| 32K | random | 0.64 | 0.0 | 0.8K | 16.7 (1.0) |
| 32K | sorted | 0.01 | 0.02 | 30.2K | 4.8 (3.5) |
[…]
| 128K | random | 0.639 | 0.0 | 0.9K | 38.0 (1.0) |
| 128K | sorted | 0.01 | 0.02 | 125K | 5.5 (6.9) |
[…]
| Seq Len | SP | GC Layer | Memory | Iter Time |
| 32K | 2 | 28 | 77G | 4.45s |
| 32K | 4 | 23 | 78G | 4.35s |
[…]
| 64K | 2 | 32 | OOM | - |
| 64K | 4 | 28 | 78G | 6.3s |
[…]
| 128K | 4 | 32 | OOM | - |
[…]
| 128K | 16 | 23 | 79G | 10.5s |
3.3 Training Strategy Analysis
Given the GPU resources and the data to be trained, we can select from various training strategies, provided that the VRAM requirements are satisfied. The main factors to consider are the degree of SP and the configuration of Gradient Checkpointing (GC), which is the number of layers where GC is enabled. If the SP degree is small, the VRAM demand is high, which forces an increase in the number of GC layers and leads to excessive additional computation. On the other hand, if the SP degree is large, although the VRAM demand is reduced, it introduces additional communication overhead. Therefore, there is a trade-off between the SP degree and GC configuration. Moreover, the optimal strategy varies for different sequence lengths. Table 4 shows that the optimal strategies for 32K, 64K, and 128K are different.
4 Hierarchical Balance Packing
In this section, we first describe how to determine optimal packing length groups (Section 4.1), then explain how each sample is assigned to a group (Section 4.2). Finally, we introduce a dynamic training pipeline for HBP (Section 4.3). The overall framework is shown in Figure 2.
4.1 Hierarchical Groups Auto-Selection
To determine the optimal packing length groups, we design a profile-based auto-selection algorithm as described in Algorithm 1. It operates in two stages: (1) finding the best training strategy for predefining a possible sequence length set (e.g., 8K, 16K, 32K, 64K, 128K) based on naive packing. (2) Deriving the final packing groups by optimizing communication overhead.
Stage 1: Find the best training strategy. The algorithm begins by initializing two empty lists, and , to store the profiling times and the corresponding strategies, respectively. For each possible input length , we compute the optimal degree of SP and the configuration of the GC with the FindBestSpCkpt function, which achieves the optimal trade-off between the degree of SP and the configuration of the GC. More details of FindBestSpCkpt are shown in Appendix I.1.
Specifically, given a packing length , we iterate all possible degrees and GC configurations and profile their iteration time. The best combination is then found by a greedy method. Once all input lengths are processed, the algorithm selects the best strategy by identifying the index that minimizes the profiling times in . The optimal strategy and input length are:
Additionally, the maximum input length and its corresponding strategy are also retained.
Stage 2: Optimize packing groups for communication.
After calculating the optimal and the corresponding , we can optimize it further to reduce communication overhead and obtain the final packing group lengths. Taking 32K packing groups in Figure 3 with SP=2 as an example, each SP process handles a split of 16K tokens. For samples longer than 16K, additional communication is required anyway. For samples shorter than 16K, we can group them into 16K packing groups, which is equivalent to training with 16K packing groups using SP=1, thereby reducing communication.
For , the smallest packing group is derived based on its optimal degree of SP , ensuring no communication overhead during sequential parallelism. Similarly, the smallest packing group for is also considered. If is smaller than , it will be merged into the range. Finally, the hierarchical packing groups is obtained. More implementation details are shown in Appendix I.
4.2 Balance Packing
After obtaining the optimal hierarchical pack groups, we distribute the dataset samples into different groups while ensuring that the metrics outlined in Section 3.1 (DBR, PR, ABR, and CR) are well-optimized, which conventional packing struggles with. We first divide the entire dataset into sub-datasets based on . For each , the following steps are executed:
(1) Packing: We pack the data in to length . This ensures low PR and DBR. Note that arbitrary packing methods are feasible.
(2) GreedyFill: We use the remaining unpacked data for a greedy fill, as larger groups struggle to fill the packed data using only their samples. A simple example illustration is shown in Appendix E.
(3) Balance Batching: Constructing balance batches based on attention complexity using heuristic solution sorting support maintaining balanced attention computation across different packing groups. Heuristic solutions are an efficient, approximate solution for large-scale datasets. More evidences are shown in Appendix D.
The procedure of balanced packing is illustrated in Algorithm 2. Since our approach inherently involves multiple levels, i.e., hierarchical packing groups, it automatically separates short- and long-context data, avoiding wasted communication overhead and imbalanced attention computation and reducing CR and ABR significantly. We also achieve low PR and DBR at multiple levels, thanks to GreedyFill. More details about the implementation are shown in Appendix J.
4.3 Dynamic Training Pipeline
Since HBP involves multi-level inputs, it is essential to design a dynamic training pipeline, enabling hot switching of different packing groups. Adaptive SP (pre-initialized multi-SP achieving zero overhead) is proposed to ensure efficient training with multi-level packing. Curriculum Learning and a Stable Loss Normalizer are proposed to improve the performance in hybrid training.
Adaptive Sequential Parallel: Each packing group is assigned an optimal training configuration , and we adopt an alternating scheme that selects the best SP degree and GC setting per packing group, as illustrated in Figure 2.
Curriculum Learning Strategy: Training on long-context tasks presents challenges because initiating training without instructional capabilities can result in significant fluctuations in the training loss, as illustrated in Figure 5. Thanks to our inherent hierarchical structure, it is straightforward to adopt a curriculum learning strategy that begins with general short-context data in the early stages of training. As training progresses, we transition to a hybrid approach that alternates training both short- and long-context data, as illustrated in Figure 2.
Stable Loss Normalizer: The training stability introduced by data packing is an important problem, as it impacts the data distribution. Previous work [6] has analyzed loss calculation, identifying two primary loss normalizers (Token-Mean) and (Sample-Mean):
where represents the local batch size within the DP group, and denotes the number of loss tokens in the batch . They argued that instability lies in the inconsistent loss of normalization. [27] also proposes sum loss without normalization to mitigate the effect of norm discrepancies. However, the sum loss introduces a trade-off: as the sequence length increases, the gradient values escalate disproportionately (1e+5), as shown in the left part of Figure 5. To address the above issues, we empirically observe (Average Token) of the global batch size across all DP groups can serve as a stable loss normalizer. The stable loss normalizer is derived from theoretical considerations, with the objective of guaranteeing that each token contributes equally to the aggregate loss. This formulation mitigates biases introduced by heterogeneous sequence lengths, varying data-parallel group sizes, and gradient accumulation strategies. We illustrate the effect of different normalization strategies under a simple setting: a data-parallel (DP) group of size , where each rank processes a local batch of samples.
Token-Mean: In the token-mean approach, the loss for each DP rank is normalized by its local token:
The global loss is then averaged across ranks:
This formulation normalizes losses rank-wise without accounting for global token distribution, potentially introducing bias when sequence lengths vary across ranks.
Sample-Mean: In the sample-mean method, losses are first normalized per sample and then averaged:
The global loss becomes:
This strategy balances individual samples but fails to reflect the total token count, giving undue weight to shorter sequences.
Stable Loss Normalizer: To eliminate such bias, we normalize by the global average token length:
Each DP’s loss is then normalized by , And the final loss is:
This normalization ensures that the contribution of each token is equal, regardless of which sample or rank it belongs to, thereby producing an unbiased global loss.
5 Experiments
5.1 Experimental Setup
Implementation Details. We use large-scale datasets: Tulu3 (32K)[14] for general tasks and LongCite (128K)[15] for long-context tasks. Both have shown strong performance across various benchmarks. They also enhance the model’s ability to handle input lengths from 0.1K to 128K tokens. We conduct experiments with the following models: LLaMA 3.1 [1], Qwen-2.5 [2], and DeepSeek-V2 (236B). Most models are trained on 32x H100 80GB GPUs using the DeepSpeed [28], while DeepSeek-V2 (236B) is trained with the Megatron-LM [29] with 256x H100 80G GPUs. In our experiments, we use DeepSpeed-Ulysses’s [10] sequence parallelism approach, and the ring-attention [11] method is also applicable. We conducted ablation experiments using the LLaMA3.1-8B model. For the Longsite dataset, approximately 2k samples are uniformly sampled. The loss normalizer for baselines without special instructions is Token-Mean, while HBP uses Ave-Token. GPU days are the evaluation metric to estimate the total training time. We use a learning rate of 1e-5, weight decay of 0.01, and adopt AdamW as our optimizer.
Evaluation. We comprehensively evaluated the LLM’s performance using OpenCompass [30]. For general tasks, several benchmark datasets were assessed, including MMLU [31], MMLU PRO [32], CMMLU [33], BBH [34], Math [35], GPQA Diamond [36], GSM8K [37], HellaSwag [38], MathBench [39], HumanEval [40], MBPP [41], IFEval [42], and Drop [43]. For long-context tasks, the evaluation including Ruler [44], NeedleBench [45], LongBench [46], and Longcite.
5.2 Main Results
| Model | General Tasks | Long Tasks | GPU Days | ||||||||
| (Type) | AVE | MMLU | BBH | IFEval | Math | GSM8k | HumanEval | Ruler (32K128K) | LongBench | LongCite | (speed up) |
| LLama3.1-8B | |||||||||||
| LongAlign-packing | 56.6 | 44.5 | 65.5 | 67.8 | 30.7 | 80.6 | 62.2 | 84.5 57.5 | 46.7 | 67.8 | 5.4 (0.97) |
| LongAlign-sorted | 57.6 | 62.7 | 65.4 | 67.8 | 32.8 | 82.2 | 61.6 | 85.8 59.9 | 46.5 | 64.0 | 33.3 (0.16) |
| ISF | 56.0 | 54.5 | 65.3 | 70.4 | 33.7 | 81.7 | 62.8 | 85.0 67.4 | 44.0 | 71.6 | 5.22 (1.0) |
[…]
| Qwen2.5-32B | |||||||||||
| ISF | 73.5 | 74.8 | 83.5 | 75.6 | 56.5 | 93.7 | 86.6 | 88.2 59.3 | 51.0 | 60.2 | 21.3 (1.0) |
[…]
| LLama3.1-70B | |||||||||||
| ISF | 72.1 | 78.9 | 83.0 | 77.6 | 44.1 | 85.3 | 76.8 | 91.8 57.1 | 50.4 | 72,7 | 44.4 (1.0) |
[…]
| DeepSeek-V2 (236B) | |||||||||||
| ISF | 71.8 | 76.8 | 84.0 | 71.1 | 41.3 | 89.0 | 78.6 | 86.6 - | 47.1 | - | 57.1 (1.0x) |
[…]
In Table 5, we compare our method HBP with LongAlign [6] and packing method ISF [8]. We also notice that LongAlign improves the average general tasks to some extent; it sacrifices performance in long tasks e.g., Ruler-128K. In contrast, our method maintains strong performance both in general short tasks and long tasks. The improvements are consistent across a wide range of model sizes, from 8B to 236B parameters. Notably, for the largest MoE model, DeepSeek-V2 (236B), our method achieves an impressive 2.4 training speed-up, reducing training time from 57.1 to 23.8 GPU days. Full results are shown in Appendix C.
| Model | Hierarchical | Balance | ABR | CR | AVE | GPU Days (speed up) |
| LLaMA3.1-8B | 0.506 | 1.0 | 56.0 | 5.22 (1.0) | ||
| LLaMA3.1-8B | ✓ | 0.288 | 0.173 | 56.4 | 4.51 (1.2) | |
[…]
| LLaMA3.1-70B | 0.506 | 1.0 | 72.2 | 44.4 (1.0) | ||
| LLaMA3.1-70B | ✓ | 0.288 | 0.173 | 72.8 | 33.3 (1.25) | |
[…]
5.3 Ablation Results
Importance of Hybrid Training. Table 10 shows that both short- and long-context data are essential for maintaining general and long-text capabilities. "Longcite (8K)" includes only sequences up to 8K. All settings use naive packing. Removing long-context data (row 1) harms long-text performance, while removing short-context data (row 2) weakens the general ability. Row 3, with partial short data, confirms these effects.
| Model | CL | AVE | LongBench |
| LLama3.1-8B-HBP | 56.6 | 41.6 | |
[…]
| LLama3.1-70B-HBP | 72.2 | 51.5 | |
[…]
| LLama3.1-8B-ISF | 56.0 | 44.0 | |
[…]
| Model | CL-Iterations | AVE | LongBench |
| LLama3.1-8B-HBP | 0 | 56.6 | 41.6 |
| LLama3.1-8B-HBP | 100 | 58.0 | 43.1 |
[…]
| LLama3.1-70B-HBP | 0 | 72.2 | 51.5 |
| LLama3.1-70B-HBP | 100 | 73.8 | 52.4 |
[…]
Components of HBP. In Table 6, we show that the Attention Balance Ratio (ABR) and Communication Ratio (CR) can be reduced significantly with hierarchical packing. In particular, ABR drops from 0.506 to 0.288, and the CR drops from 1.0 to 0.173. By balancing the batching data with a similar complexity of attention computation, we have much more balanced mini-batches with low ABR, from 0.288 to 0.002. Overall, we achieve a 1.4 speedup. Similar results can be observed in the larger LLaMA-3.1-70B model.
Curriculum Learning. Table 8 presents the impact of curriculum learning on HBP. Starting with short tasks and gradually mixing in long-context tasks benefits training and convergence. We applied a similar curriculum strategy to the naive ISF baseline using advanced sampling, which also showed improvements. This confirms the general effectiveness of curriculum learning for long-context SFT. Notably, HBP naturally separates short and long contexts, making curriculum learning easier to apply.
Iterations for Curriculum Learning training. Table 8 presents detailed ablation studies across different models. As shown, even a simple curriculum learning setup, adding 100 steps of early-stage training with short samples, already yields noticeable improvements. To ensure training stability, we ultimately adopt a setting with 500 steps of short-sample-only training in our final experiments.
| Dataset | Pack Len | AVE | LongBench | Ruler-128K |
| Tulu3 | 32K | 57.5 | 43.0 | 52.5 |
| Longcite | 128K | 18.8 | 16.7 | 68.5 |
| Tulu3 + Longcite(8K) | 32K | 58.6 | 43.2 | 62.1 |
[…]
| Loss Normalizer | AVE | LongBench | Ruler-128K | Longcite |
| Sum | 56.7 | 42.5 | 65.2 | 70.5 |
| Sample-Mean | 55.5 | 42.9 | 46.1 | 70.3 |
| Token-Mean | 56.6 | 41.6 | 67.5 | 70.6 |
[…]
Stable Loss Normalizer. We compared several loss normalization methods by training models using different normalizers while keeping all other training configurations consistent. As shown in Table 10, the Ave-Token loss normalizer achieved the highest performance in both general tasks, Average (AVE) 57.6, and long context tasks LongBench 43.1, Ruler-128K 70.8, and LongSite 71.2.
Importance of Hierarchical Groups Auto-Selection. Table 11 compares manual (row 2) and automatic (row 3) packing groups. Automatic selection of groups (based on Appendix F) achieves greater speedup, reducing training time from 4.25 to 3.73 GPU days.
| Model | Groups | Selection | SP | ABR | CR | AVE | LongBench | GPU Days (speed up) |
| LLama3.1-8B-ISF | [128K] | manual | [8] | 0.506 | 1.0 | 56.0 | 44.0 | 5.22 (1.0) |
| LLama3.1-8B-HBP | [32K, 128K] | manual | [2, 8] | 0.002 | 1.0 | 57.9 | 43.2 | 4.25 (1.22) |
[…]
Compared with FlexSP. As shown in Table 12, HBP achieves lower ABR, DBR, and PR than FlexSP, while maintaining a similar CR, contributing to faster training (4.35 to 3.73 GPU days), even considering only the model computation time. Furthermore, FlexSP introduces a per-batch grouping overhead (5-15 seconds), which further slows down training (see Appendix B). HBP performs global data-level operations in advance with negligible data overhead.
Negligible Overhead. The overhead of HBP is negligible: Auto-Selection takes minutes and Packing seconds, together contributing less than of training time (see Appendix G).
| Model | SP | CL | PR | DBR | CR | ABR | AVE | LongBench | GPU Days (speed up) |
| LLama3.1-8B-ISF | [8] | 0 | 0.001 | 1.0 | 0.506 | 56.0 | 44.0 | 5.22 (1.0) | |
| LLama3.1-8B-FlexSP [24] | [1,2,4,8] | 0.04 | 0.064 | 0.171 | 0.36 | 56.1 | 41.8 | 4.35 (1.2) | |
| LLama3.1-8B-HBP | [1,8] | 0 | 0.001 | 0.173 | 0.002 | 56.6 | 41.6 | 3.73 (1.4) | |
[…]
6 Conclusion and Limitations
In this paper, we proposed Hierarchical Balance Packing (HBP), a novel strategy to address workload imbalances in long-context LLM training through multi-level data packing and a dynamic training pipeline. Due to limitations in computational resources, we have not conducted experiments on longer contexts, such as 256K or 512K. Additionally, we have not validated HBP on other post-training tasks, such as RLHF or DPO. We leave these explorations to future work.
Acknowledgements
This work was supported by the National Key Research and Development Program of China (Grant No. 2023YFB4405102), the Postdoctoral Fellowship Program and the China Postdoctoral Science Foundation (Grant No. BX20250487), and in part by the Natural Science Foundation of Hunan Province (Grant No. 2024JJ6525).
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the contributions and scope of the paper?
Answer: [Yes]
Justification: The main claims in the abstract and introduction accurately reflect the paper’s contributions and scope.
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: The paper provides a brief discussion of the limitations in the Conclusion section.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory assumptions and proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Justification: The paper is primarily based on empirical observations and proposes heuristic approaches, which are not fully supported by formal mathematical proofs or a complete set of theoretical assumptions.
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental result reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Justification: The experimental results can be fully reproduced based on the information provided in the paper, and the authors have indicated that the code will be released in the future.
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [N/A]
Justification: The code and data are not yet publicly available, and although the authors plan to release them later, they are currently not provided with sufficient instructions in the supplemental material to ensure faithful reproduction of the results.
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental setting/details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Justification: The full details are provided in the main text and supplemental material.
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment statistical significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [Yes]
Justification: The results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments compute resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: The paper provides sufficient information about the computational resources used, including details in both the main text and the Appendix.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code of ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: The research presented in the paper fully conforms to the NeurIPS Code of Ethics in all respects.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [N/A]
Justification: There is no societal impact of the work performed.
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [N/A]
Justification: The paper poses no such risks.
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: All resources used in the paper comply with their respective licenses and terms of use.
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [N/A]
Justification: The paper does not release new assets.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and research with human subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: The paper does not involve crowdsourcing nor research with human subjects.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional review board (IRB) approvals or equivalent for research with human subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: The paper does not involve crowdsourcing nor research with human subjects.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
16.
Declaration of LLM usage
Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
Answer: [N/A]
Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
Guidelines:
The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Packing Strategy Results
Table 13 illustrates the complexity of different packing strategies and their corresponding final training times. It also shows that the training indicators Data Balance Ratio (DBR), Padding Ratio (PR), and Attention Balance Ratio (ABR) are strongly correlated with the training time, emphasizing the effectiveness of these indicators. Based on the computational complexity of packing strategies and their training time, and accuracy performance, we ultimately selected ISF as our naive packing baseline.
| Packing Strategy | Complexity | DBR | ABR | PR | Ave | LongBench | GPU Days (speed up) |
| No | - | 0.63 | 0.72 | 0 | 56.5 | 41.8 | 38.3 (1.00) |
| Random | O(N) | 0 | 0.53 | 0.03 | 56.2 | 42.0 | 5.57 (6.87) |
[…]
| FFS | O(NM) | 0 | 0.508 | 0.01 | 56.3 | 42.0 | 5.24 (7.30) |
| FFD | O(NM) | 0 | 0.515 | 0.01 | 56.4 | 42.4 | 5.48 (6.98) |
| BFS | O(NM) | 0 | 0.508 | 0.01 | 56.8 | 41.7 | 5.24 (7.30) |
| SPFHP | O(N+S2) | 0 | 0.512 | 0.01 | 56.2 | 42.3 | 5.40 (7.10) |
Appendix B Results of FlexSP data grouping
The experiments were conducted with a sequence length of 128K using 32 GPUs. The results for FlexSP were obtained by directly testing with the officially released code. The FlexSP data group method incurs a non-negligible overhead of 5–15 seconds per batch, which cannot always be hidden by the training time shown in Table 14 Rows 3 and 4.
| Model | Dataset | Group method | Data Time | Model Time | Total Time |
| GPT-7b | github | FlexSP | 0.25s | 15.4s | 15.6s |
| GPT-7b | CommonCrawl | FlexSP | 0.3s | 13.0s | 13.3s |
| GPT-7b | Wikipedia | FlexSP | 4.41s | 2.4s | 6.81s |
| GPT-7b | Tulu3 + Longcite | FlexSP | 3.43s | 3.02s | 6.45s |
[…]
Appendix C Full Results
Tables 15 and 16 present our complete results for general tasks and long-context tasks, respectively. These results collectively validate the effectiveness of our HBP method. The other dense models all use the GQA architecture, while Deepseek-V2 adopts the MLA structure, which involves more communication due to a larger number of K and V heads, similar to MHA. HBP effectively reduces the communication overhead, making the speedup more significant.
| Model | MMLU | BBH | IFEval | Math | GSM8k | HumanEval | mmlu_pro | cmmlu | GPQA | Drop | MBPP | hellaswag | mathbench-a | mathbench-t | AVE | GPU Days (speed up) |
| LLama3.1-8B | ||||||||||||||||
| LongAlign-packing | 44.5 | 65.5 | 67.8 | 30.7 | 80.6 | 62.2 | 26.5 | 48.4 | 28.8 | 71.5 | 63.8 | 80.5 | 45.6 | 75.6 | 56.6 | 7.4 (0.7) |
| LongAlign-sorted | 62.7 | 65.4 | 67.8 | 32.8 | 82.2 | 61.6 | 38.3 | 43.6 | 29.3 | 65.1 | 63.0 | 69.6 | 50.3 | 73.9 | 57.6 | 33.3 (0.16) |
| ISF | 54.5 | 65.3 | 70.4 | 33.7 | 81.7 | 62.8 | 34.6 | 39.8 | 24.2 | 65.4 | 61.8 | 67.9 | 46.7 | 74.7 | 56.0 | 5.22 (1.0) |
[…]
| Qwen2.5-32B | ||||||||||||||||
| ISF | 74.8 | 83.5 | 75.6 | 56.5 | 93.7 | 86.6 | 59.6 | 77.6 | 37.9 | 82.7 | 80.1 | 93.5 | 64.1 | 63.5 | 73.5 | 21.33 (1.0) |
[…]
| LLaMA3.1-70B | ||||||||||||||||
| ISF | 78.9 | 83.0 | 77.6 | 44.1 | 85.3 | 76.8 | 55.0 | 66.9 | 41.4 | 81.9 | 76.6 | 89.3 | 65.1 | 88.0 | 72.1 | 44.40 (1.0) |
[…]
| Qwen2.5-72B | ||||||||||||||||
| ISF | 83.9 | 86.0 | 79.3 | 57.2 | 94.6 | 85.9 | 66.2 | 84.7 | 45.4 | 84.6 | 84.8 | 93.7 | 74.1 | 95.0 | 79.6 | 47.10 (1.0) |
[…]
| DeepSeek-V2 (236B) | ||||||||||||||||
| ISF | 76.8 | 84.0 | 71.1 | 41.3 | 89.0 | 78.6 | 54.2 | 76.9 | 37.9 | 77.2 | 76.7 | 90.2 | 68.4 | 92.3 | 71.8 | 57.10 (1.0) |
[…]
| Model | Ruler (32K | 128K) | NeedleBench (32K | 128K) | LongBench | Longcite | GPU Days (speed up) |
| LLama3.1-8B | |||||
| LongAlign-packing | 84.5 | 57.5 | 87.9 | 85.0 | 46.7 | 67.8 | 7.4 (0.7) |
| LongAlign-sorted | 85.6 | 60.0 | 92.4 | 88.9 | 46.6 | 64.0 | 33.3 (0.16) |
| ISF | 85.0 | 67.4 | 92.1 | 90.1 | 44.5 | 71.6 | 5.22 (1.0) |
[…]
| Qwen2.5-32B | |||||
| ISF | 88.2 | 59.3 | 94.5 | 84.6 | 51.0 | 60.2 | 21.33 (1.0) |
[…]
| LLaMA3.1-70B | |||||
| ISF | 91.8 | 57.1 | 95.5 | 92.6 | 50.4 | 72.7 | 44.4 (1.0) |
[…]
| Qwen2.5-72B | |||||
| ISF | 92.6 | 58.5 | 94.5 | 90.8 | 50.0 | 64.3 | 47.1 (1.0) |
[…]
| DeepSeek-V2 (236B) | |||||
| ISF | 86.6 | - | 96.0 | - | 47.1 | - | 57.1 (1.0) |
[…]
Appendix D More Details of Attention Imbalance Problem
Imbalance of attention complexity: As deduced from Megatron-LM paper, gives a packed sequence containing n sub-sequences with total lengths, each transformer layer computes cost can be approximated as:, where is the length of the i-th sub-sequence, B is the batch size, and h is the hidden size. Therefore, variance in attention complexity directly impacts compute cost.
Effectiveness of Balance Batching:
Sorting is indeed a heuristic solution, and it cannot be strictly proven mathematically to guarantee perfect balance. However, based on the results across different datasets, it achieves near-perfect balance (low ABR approaching 0).
Heuristic solutions are a common and efficient approximate solution for solving problems on large-scale datasets. Our method ensures that iterations maintain attention to balance across different datasets over 99.6% of training shown in Table 17, demonstrating the strong generalization capability of our method.
Balance Evidence: The left side in Figure 6 shows baseline packing with high attention cost variance across DP groups, while the right side shows our sorting-based method, which significantly reduces this variance and improves efficiency.
| Dataset | Packing method | ABR |
| Tulu3 | HBP | 0.001 |
| OpenHermes | HBP | 0.001 |
| Tulu3 + Longcite | HBP | 0.002 |
| OpenHermes + Longcite | HBP | 0.004 |
| Tulu3 + Longwriter | HBP | 0.003 |
Appendix E More Details of GreedyFill.
GreedyFill aims to reduce intra-pack padding tokens and improve Ave-T. It operates within local packing groups by selecting a small number of sequences from other shorter packing groups to efficiently fill residual space, as shown in Figure 7.
Appendix F More Details of Hierarchical Groups Auto-Selection.
The necessity of auto-selection: Table 19 presents an example of 32K token-length training, showcasing various SP degrees and GC configurations. The second row highlights the minimal configuration that satisfies memory constraints. While this configuration adheres to memory requirements, it fails to achieve optimal performance. In contrast, the fourth row illustrates a more balanced and effective configuration with , achieving the fastest speed. These experiments demonstrate that, given a specific packing length, there are significant performance differences among various SP degrees and GC configurations.
The final selection groups result: Table 19 presents the training speed of the optimal SP degrees and GC configurations for various sequence lengths . The second row of the table shows that our optimal configuration is groups = 16K, with the corresponding training strategy being . This indicates that the optimal training strategy is distinct for different packing groups. The evidence above emphasizes the importance of searching for the best groups and their corresponding training strategies, i.e., Auto-Group Selection.
| SP | GC Layer | AVE | Memory | Iter Time | GPU Days |
| 1 | 32 | - | OOM | - | - |
| 2 | 28 | 57.5 | 78G | 3.01 s | 3.73 |
| 4 | 23 | 57.8 | 78G | 2.95 s | 3.37 |
[…]
| 16 | 0 | 56.6 | 76G | 3.60 s | 3.93 |
| Packing Groups | SP | GC Layer | Memory | Iter Time |
| 8k | 2 | 8 | 77 G | 2.69 s |
[…]
| 32k | 8 | 8 | 76 G | 2.83 s |
| 64k | 4 | 28 | 78 G | 3.01 s |
| 128k | 8 | 28 | 79 G | 3.05 s |
Appendix G Overhead Analysis
Balance Packing and Auto-Selection overhead relative to training time. Both steps together remain consistently negligible: Auto-Selection requires at most minutes and Packing only seconds, while training spans from to minutes. Overall, the combined overhead is below of training time, confirming that preprocessing costs are insignificant compared to model training.
| Model | Dataset | Auto-Selection Time | Balance Packing Time | Training Time | Overhead Ratio |
| LLaMA3.1-8B | Tulu3 + Longcite | 3.25 min | 5 sec | 168 min | 2% |
| LLaMA3.1-70B | Tulu3 + Longcite | 15 min | 5 sec | 1400 min | 1% |
G.1 Overhead of Profiling
Memory Profiling.
The memory profiling cost depends on the sequence length set (Seq) processed by the device:
Time Profiling.
The total number of search strategies () is determined by the packing length set () and the Sequence Parallel () settings. profile_iter can be adjusted based on actual iteration times, with typical values ranging from to :
Total Cost (Example).
For instance, with , and , , for LLaMA3.1-8B:
Considering the 3000+ training iterations, totaling iterations (2%) is negligible compared to the overall training time.
G.2 Overhead of Balance Packing
For our training dataset of million samples, the overhead introduced by Balance Packing is consistently within seconds.
Complexity.
The computational complexity of Balance Packing can be expressed as:
where is the number of samples, is the number of packing groups, is the number of ISF iterations, and denotes the selected packing set.
Since , , and , the overall complexity reduces to a sublinear class relative to . Hence, scaling to substantially larger datasets incurs only minimal additional overhead.
G.3 Scalability of the Proposed Method
Large-scale dataset.
Our profiling is performed for only a small number of iterations (). As the dataset size increases, the overall overhead ratio () becomes even lower. Meanwhile, the complexity of Balance Packing is , so the additional cost introduced is negligible as the dataset scales.
Appendix H Dataset Generalization
H.1 OpenHermes
We also provide some of our experimental results on OpenHermes. For example, Table 21 shows that the results are consistent with Tulu3 under different packing strategies. The label 22 shows that our HBP is also effective in a 128K training setting.
| Packing Strategy | Complexity | DBR | ABR | PR | Ave | LongBench | GPU Days (speed up) |
| - | - | 0 | 0.878 | 0.676 | 48.2 | 46.4 | 11.4 (1.0) |
| random | O(N) | 0 | 0.648 | 0.089 | 51.44 | 47.8 | 3.64 (3.1) |
[…]
| FFS | O(NM) | 0 | 0.638 | 0.022 | 52.4 | 48.4 | 3.29 (3.5) |
| FFD | O(NM) | 0 | 0.71 | 0.022 | 50.6 | 48.4 | 3.45 (3.3) |
| BFS | O(NM) | 0 | 0.64 | 0.022 | 52.3 | 47.7 | 3.33 (3.4) |
| SPFHP | O(N+S2) | 0 | 0.71 | 0.029 | 52.0 | 47.8 | 3.47 (3.3) |
| Hierarchical | Balance | ABR | AVE | LongBench | Ruler-32K | Longsite | GPU Days (speed up) |
| 0.513 | 52.1 | 47.2 | 87.7 | 72.1 | 2.95 (1.0) | ||
| ✓ | 0.269 | 53 | 47.1 | 89.1 | 72.0 | 2.33 (1.27) | |
[…]
H.2 LongWriter
Figure 8 and Table 23 present the results of our HBP hybrid training on Tulu3 and LongWriter. The x-axis represents the user instruction required length, and the y-axis represents the model output length based on the Long-Writer paper. A stronger linear correlation indicates better instruction-following capability of the model. The left side of the Figure shows the original baseline results without using the Long-Writer dataset. The middle shows the results of naive packing, and the right shows the results of HBP. The experiment demonstrates that our method is equally effective, achieving consistent acceleration across both models.
| Model | Dataset | AVe | LongBench | LongWriter | GPU Days (speed up) |
| Llama3.1-8B-ISF | Tulu3 + LongWriter | 57.3 | 40.9 | 67.0 | 4.4 (1.0) |
[…]
Appendix I Implementation Details of Hierarchical Groups Auto-Selection
I.1 FindBestSpCkpt
Algorithm 3 determines the optimal gradient checkpointing strategy by evaluating all possible strategies.
Initialization: Start with empty lists and for profiling times and configurations, respectively.
Iterate Over Strategies: For each strategy :
Compute the best gradient checkpointing configuration () using the GreedyProfileCkpt function.
Use ProfileTime to profile the execution time for the model with the given configuration and append it to .
Find Optimal Strategy: Identify the index of the minimum profiling time in using .
Return Best Configuration: Return the SP degree and corresponding gradient checkpointing configuration .
I.2 GreedyProfileCkpt
The algorithm 4 estimates the number of gradient checkpoint layers required for a given strategy and . Here, we provide a more detailed explanation. We use the "pynvml" library to monitor GPU memory usage for memory analysis.
Memory Profiling:
Step 1: Given an input length L, enable Gradient Checkpointing for all layers (32 for 7B), and record the remaining GPU memory as m1.
Step 2: Given the same input length L, enable Gradient Checkpointing for only a subset of layers (e.g., 20 layers), and record the remaining GPU memory as m2.
Step 3: From the above, we can estimate the average memory saved per layer with Gradient Checkpointing: ave_m = (m1 - m2) / (32 - 20)
By running only 10 iterations, we can obtain a reliable estimate of the memory saved per layer when using Gradient Checkpointing.
Memory And Time Consumption Report:
For a given input length L, we set different sequence parallelism (SP) configurations. Based on the remaining GPU memory and the previously estimated ave_m, we determine the number of layers to apply Gradient Checkpointing. Then, we run 10 iterations to record the corresponding memory usage and iteration time.
Full workflow:
Initialization: Obtain the configurations using the minimum and maximum number of gradient checkpointing layers (based on empirical observations):
Memory Profiling: Profile the remaining memory for () and ().
Memory Slope Calculation: Compute the average memory slope () as:
Checkpointing Layer Estimation: Estimate the required number of checkpoints () as:
Appendix J Implementation Details of Balance Packing
J.1 GroupData
Given a data set and predefined hierarchical lengths , we evaluate the length of each data set to determine the interval to which it belongs, assigning it to the corresponding group . The detailed procedure is outlined in Algorithm 5.
J.2 GreedyFill
For a given packing group with the corresponding length , we iterate through the smaller dataset partitions and greedily fill the group within the current . The detailed procedure is illustrated in Algorithm 6.
J.3 Attention Balance Sort Function
First, we calculate the attention complexity for all data within the given packing group . Then, we sort the elements in based on their attention complexity and construct mini-batches according to the global token number requirements as shown in Algorithm 7.